a uniform plank PQ of length 10m and mass (m)kg rests on 2 supports A and B, where /PA/=/BQ/=1m. A load of mass 8kg is placed on the plank at point C such that /AC/ = 3.5m. If the reaction is 100N, calculate the (i) value of m (ii)reaction at A(take g=10ms^-2)
1
Expert's answer
2020-09-16T10:10:31-0400

Make a drawing:



Since the plank does not move, all forces and torques are balanced. Forces:


"R_A+R_B-m_Cg-m_pg=0,\\"


torques (consider A as the pivot):


"\frac{m_p}{10}g\cdot0.5-\frac{9}{10}m_pg\cdot4.5-m_Cg\cdot3.5+100\cdot8=0,"

the solution is


"m_p=13\text{ kg},\\
R_A=110\text{ N}."